# On deviation probabilities in non-parametric regression

### v2

@@ Abstract.
This paper is devoted to the problem of determining the concentration bounds that are achievable in non-parametric regression. We consider the setting where features are supported on a bounded subset of $\mathbb{R}^{d}$, the regression function is Lipschitz, and the noise is only assumed to have a finite second moment. We first specify the fundamental limits of the problem by establishing a general lower bound on deviation probabilities, and then construct explicit estimators that achieve this bound. These estimators are obtained by applying the median-of-means principle to classical local averaging rules in non-parametric regression, including nearest neighbors and kernel procedures.
@@ 1.1. Setting and main results
Let $(X,Y)$ be a pair of random variables, where $X$ has distribution $\mu$ on $\mathbb{R}^{d}$, for $d\geq 1$, and $Y$ is real-valued and satisfies $\mathbb{E}[Y^{2}]<\infty$. The regression function is defined for $\mu$-almost every $x\in\mathbb{R}^{d}$ as
@@
where the infimum is taken over all measurable functions $g:\mathbb{R}^{d}\to\mathbb{R}$ such that $\mathbb{E}[g(X)^{2}]<\infty$. In other words, the regression function is an optimal approximation of $Y$ by a square-integrable function of $X$, with respect to the $L_{2}$ risk.
When the distribution of the pair $(X,Y)$ is unknown, one cannot predict $Y$ using $r(X)$. However, assuming that one has access to an i.i.d. sample
@@
with the same distribution as $(X,Y)$, then one can use the data ${\mathcal{D}}_{n}$ in order to construct an estimate $\widehat{r}_{n}$ of the function $r$.
@@
The class ${\mathcal{F}}={\mathcal{F}}_{\rho,\sigma}$, with $\rho,\sigma>0$, is the class of distributions $(X,Y)$ satisfying:
@@
up to some constants depending on $d$ only, see [36] and [13], Chapter 3. In particular, this rate is achieved by nearest neighbors or kernel procedures. Let us note that usually, this minimax rate is written with $D^{-d}$ instead of $\rho$, where $D$ is the diameter of $S$. However, those two quantities are clearly related by the inequality $\rho\leq D^{-d}$, which comes from (1) applied to $\varepsilon=D$. Moreover, one may check that lower bounds for the $L_{2}$ risk are actually obtained in situations where equation (1) is satisfied, namely with $X$ uniform on a $d$-dimensional hypercube.
@@
For any $\delta\in]0,1/16]$, for any $\rho>0$ and $\sigma>0$, for any regression estimate $\widehat{r}_{n}$, there exists a distribution $(X,Y)\in{\mathcal{F}}_{\rho,\sigma}$ as defined in Assumption 1 such that, when $X\sim\mu$ independent of ${\mathcal{D}}_{n}$, we have
$\mathbb{P}\left(\left|\widehat{r}_{n}(X)-r(X)\right|\geq a_{d}\left(\frac{\sigma^{2}}{\rho n}\ln\left(\frac{1}{16\delta}\right)\right)^{\frac{1}{d+2}}\right)\leq\delta\,,$
@@
For a variety of base estimates, including nearest neighbors and kernel estimates, we show that, when $(X,Y)\in{\mathcal{F}}_{\rho,\sigma}$, and when $\sigma$ and $\rho$ are known, the median-of-means estimate satisfies the following concentration inequality: for all $\delta\in[e^{-nc_{\mathcal{F}}+1},1[$, and for all $x\in S$, when the number of blocks $m$ is chosen as $m=\left\lceil\ln(1/\delta)\right\rceil$, we have, for all $x\in S$,
@@
which, in view of Theorem 1, is the optimal concentration bound.
@@ 1.2. Related work
To our knowledge, there are only very results on deviation probabilities in non-parametric regression, at the exception of [18] for the $k$-nn estimate. However, in the latter, the noise is assumed sub-Gaussian. Therefore, the first contribution of the present paper is to delineate the fundamental limits of the problem by providing a general lower bound when only a second moment assumption on the noise is made (see Theorem 1). Our second contribution is to show that this bound can in fact be achieved by combining local averaging rules and the MoM principle.
@@
Finally, let us note that the minimax deviation inequalities (2) are obtained by optimizing on the tuning parameter of the base estimate (e.g., the number $k$ of neighbors for $k$-nn estimates, the bandwidth $h$ for kernel estimates), and this optimization step typically requires the knowledge of $\rho$ and $\sigma$. In this respect, an open question would be to design procedures to choose this tuning parameter in an adaptive way. For the $L_{2}$ risk, this is possible for instance by splitting the sample or cross-validation, as explained for example in [13], Chapters 7 and 8. Unfortunately, this issue seems to be more complicated in the present case of concentration bounds.
@@ 1.3. Organization of the paper
Section 2 is devoted to the proof of Theorem 1. Then Section 3 provides a guideline for proving  (2). It also contains two general observations about our estimators: the fact that they are robust to the presence of outliers in the sample, and the fact that they can be turned into $\delta$-independent estimators, using a strategy introduced by [11]. In the next two sections, Equation (2) is established for various choices of local averaging procedures, namely nearest neighbor methods in Section 4, and kernel and partitioning methods in Section 5. Let us point out that, for partitioning estimates (Section 5.2), we are able to obtain a uniform control on $\sup_{x\in S}|\widehat{r}_{n}^{\mathsf{mom}}(x)-r(x)|$. Finally, Section 6 gathers the proofs of several technical results.
@@ 2. Lower bound
where $X\sim\mu=\mathrm{Unif}(S)$, where the noise $\varepsilon\sim{\mathcal{N}}(0,\sigma^{2})$ is independent of $X$, and where $r$ is a $1$-Lipschitz function on $S$. Notice that this model satisfies Assumption 1, meaning that $(X,Y)\in\mathcal{F}$. Indeed, the support $S$ is bounded with diameter $D=A\sqrt{d}$ and, for all $x\in S$ and $\varepsilon\in(0,D]$, one has
@@ 3.2. Robustness
In addition to exhibiting concentration properties, the estimate $\widehat{r}^{\mathsf{mom}}_{n}$ also stands out through its strong robustness to outliers. In this section, we consider the contamination scheme introduced by [20]. We assume that the index set $\{1,\dots,n\}$ is divided into two disjoint subsets: the subset ${\mathcal{I}}$ of inliers, and the subset ${\mathcal{O}}$ of outliers. The sequence $(X_{i},Y_{i})_{i\in{\mathcal{I}}}$ is i.i.d. with the same law as $(X,Y)\in{\mathcal{F}}$. No assumption is made on the variables $(X_{i},Y_{i})_{i\in{\mathcal{O}}}$. We denote by $\mathbb{P}_{{\mathcal{O}}\cup{\mathcal{I}}}$ the distribution corresponding to such a contaminated sample.
@@
More generally, define $\varepsilon$ as the proportion of outliers in the original sample and assume that $m>2|{\mathcal{O}}|=2\varepsilon n$, then a straightforward adaptation of the previous reasoning shows that
$\mathbb{P}\left(\sum_{j=1}^{m}{\mathbf{1}}_{\left\{\left|\widehat{r}^{(j)}(x)-r(x)\right|\geq t+s\right\}}\geq\frac{m}{2}-\varepsilon n\right)\leq 2^{m}p_{t+s}(x)^{\frac{m}{2}-\varepsilon n}\,,$
which is less than $e^{-m}$ for $p_{t+s}(x)\leq(2e)^{-\frac{2m}{m-2\varepsilon n}}$.
@@ 4. Nearest neighbors estimation
We first propose to illustrate this result on two specific examples of nearest neighbors rules: the uniform $k$-nearest neighbors estimate (Subsection 4.1) and the bagged 1-nearest neighbor estimate (Section 4.2). As we will see, both satisfy the concentration inequality (2).
@@ 4.2. Bagging and Nearest Neighbors
Here we focus on sampling with replacement and get the following bound, whose proof can be found in Section 6.3.

### v3

@@ Abstract.
This paper is devoted to the problem of determining the deviation bounds that are achievable in non-parametric regression. We consider the setting where features are supported on a bounded subset of $\mathbb{R}^{d}$, the regression function is Lipschitz, and the noise is only assumed to have a finite second moment. We first specify the fundamental limits of the problem by establishing a general lower bound on deviation probabilities, and then construct explicit estimators that achieve this bound. These estimators are obtained by applying the median-of-means principle to classical local averaging rules in non-parametric regression, including nearest neighbors and kernel procedures.
Index Terms: Non-parametric regression, Median of Means, Sub-Gaussian estimators.
Mathematics Subject Classification (2020): 62G08, 62G15, 62G35.
@@ 1.1. Setting and main results
Let $(X,Y)$ be a pair of random variables, where $X$ has distribution $\mu$ on $\mathbb{R}^{d}$, for $d\geq 1$, and $Y$ is real-valued and satisfies $\mathbb{E}[Y^{2}]<\infty$. We denote by $Q$ the distribution of the pair $(X,Y)$. The regression function is defined for $\mu$-almost every $x\in\mathbb{R}^{d}$ as
@@
where the infimum is taken over all measurable functions $g:\mathbb{R}^{d}\to\mathbb{R}$ such that $\mathbb{E}[g(X)^{2}]<\infty$. In other words, the random variable $g(X)$ is an optimal approximation of $Y$ by a square-integrable function of $X$, with respect to the $L_{2}$ risk.
When the distribution $Q$ of the pair $(X,Y)$ is unknown, one cannot predict $Y$ using $r(X)$. However, assuming that one has access to an i.i.d. sample
@@
with the same distribution as $(X,Y)$, one can use the data ${\mathcal{D}}_{n}$ in order to construct an estimate $\widehat{r}_{n}$ of the function $r$.
@@
The class ${\mathcal{F}}={\mathcal{F}}_{\rho,\sigma}$, with $\rho,\sigma>0$, is the class of distributions $Q$ for $(X,Y)$ satisfying:
@@
up to some constants depending on $d$ only, see[14], Theorem 3.2, and [39]. In particular, this rate is achieved by nearest neighbors or kernel procedures. Let us note that usually, this minimax rate is written with $D^{-d}$ instead of $\rho$, where $D$ is the diameter of $S$. However, those two quantities are clearly related by the inequality $\rho\leq D^{-d}$, which comes from (1) applied to $\varepsilon=D$. Moreover, one may check that lower bounds for the $L_{2}$ risk are actually obtained in situations where equation (1) is satisfied, namely with $X$ uniform on a $d$-dimensional hypercube.
@@
For any $\delta\in]0,1/16]$, for any $\rho>0$ and $\sigma>0$, for any regression estimate $\widehat{r}_{n}$, there exists a distribution $Q\in{\mathcal{F}}_{\rho,\sigma}$ as defined in Assumption 1 such that, when $X\sim\mu$ independent of ${\mathcal{D}}_{n}$, we have
$\mathbb{P}\left(\left|\widehat{r}_{n}(X)-r(X)\right|\geq a_{d}\left(\frac{\sigma^{2}}{\rho n}\ln\left(\frac{1}{16\delta}\right)\right)^{\frac{1}{d+2}}\right)\geq\delta\,,$
@@
For a variety of base estimates, including nearest neighbors and kernel estimates, we show that, when $Q\in{\mathcal{F}}_{\rho,\sigma}$, and when $\sigma$ and $\rho$ are known, the median-of-means estimate satisfies the following deviation inequality: for all $\delta\in[e^{-nc_{\mathcal{F}}+1},1[$, and for all $x\in S$, when the number of blocks $m$ is chosen as $m=\left\lceil\ln(1/\delta)\right\rceil$, we have, for all $x\in S$,
@@
which, in view of Theorem 1, is the optimal deviation bound.
@@ 1.2. Related work
To our knowledge, there are only very few results on deviation probabilities in non-parametric regression, at the exception of [19] for the $k$-nn estimate. However, in the latter, the noise is assumed sub-Gaussian. Therefore, the first contribution of the present paper is to delineate the fundamental limits of the problem by providing a general lower bound when only a second moment assumption on the noise is made (see Theorem 1). Our second contribution is to show that this bound can in fact be achieved by combining local averaging rules and the MoM principle.
@@
Finally, let us note that the minimax deviation inequalities (2) are obtained by optimizing on the tuning parameter of the base estimate (e.g., the number $k$ of neighbors for $k$-nn estimates, the bandwidth $h$ for kernel estimates), and this optimization step typically requires the knowledge of $\rho$ and $\sigma$. In this respect, an open question would be to design procedures to choose this tuning parameter in an adaptive way. For the $L_{2}$ risk, this is possible for instance by splitting the sample or cross-validation, as explained for example in [14], Chapters 7 and 8. Unfortunately, this issue seems to be more complicated in the present case of deviation bounds.
@@ 1.3. Organization of the paper
In Section 2, we give the first steps for the proof of Theorem 1. Then Section 3 provides a guideline for proving  (2). It also contains two general observations about our estimators: the fact that they are robust to the presence of outliers in the sample, and the fact that they can be turned into $\delta$-independent estimators, using a strategy introduced by [11]. In the next two sections, Equation (2) is established for various choices of local averaging procedures, namely nearest neighbor methods in Section 4, and kernel and partitioning methods in Section 5. Let us point out that, for partitioning estimates (Section 5.2), we are able to obtain a uniform control on $\sup_{x\in S}|\widehat{r}_{n}^{\mathsf{mom}}(x)-r(x)|$. Finally, Section 6 gathers the proofs of several technical results.
@@ 2. Lower bound
where $X\sim\mu=\mathrm{Unif}(S)$, where the noise $\varepsilon\sim{\mathcal{N}}(0,\sigma^{2})$ is independent of $X$, and where $r$ is a $1$-Lipschitz function on $S$. Notice that this model satisfies Assumption 1, meaning that $Q\in\mathcal{F}$. Indeed, the support $S$ is bounded with diameter $D=A\sqrt{d}$ and, for all $x\in S$ and $\varepsilon\in(0,D]$, one has
@@ 3.2. Robustness
In addition to exhibiting concentration properties, the estimate $\widehat{r}^{\mathsf{mom}}_{n}$ also stands out through its strong robustness to outliers. In this section, we consider a general contamination scheme which was introduced by [12], and which was also considered in [21]. We assume that the index set $\{1,\dots,n\}$ is divided into two disjoint subsets: the subset ${\mathcal{I}}$ of inliers, and the subset ${\mathcal{O}}$ of outliers. The sequence $(X_{i},Y_{i})_{i\in{\mathcal{I}}}$ is i.i.d. with the same law as $(X,Y)\sim Q\in{\mathcal{F}}$. No assumption is made on the variables $(X_{i},Y_{i})_{i\in{\mathcal{O}}}$. We denote by $\mathbb{P}_{{\mathcal{O}}\cup{\mathcal{I}}}$ the distribution corresponding to such a contaminated sample.
@@
More generally, define $\alpha$ as the proportion of outliers in the original sample and assume that $m>2|{\mathcal{O}}|=2\alpha n$, then a straightforward adaptation of the previous reasoning shows that
$\mathbb{P}\left(\sum_{j=1}^{m}{\mathbf{1}}_{\left\{\left|\widehat{r}^{(j)}(x)-r(x)\right|\geq t+s\right\}}\geq\frac{m}{2}-\alpha n\right)\leq 2^{m}p_{t+s}(x)^{\frac{m}{2}-\alpha n}\,,$
which is less than $e^{-m}$ for $p_{t+s}(x)\leq(2e)^{-\frac{2m}{m-2\alpha n}}$. Results in the same spirit are available in the literature for the problems of mean estimation [33], Lemma 5.4, and robust regression  [35], Lemma D.1.
@@ 4. Nearest neighbors estimation
We first propose to illustrate this result on two specific examples of nearest neighbors rules: the uniform $k$-nearest neighbors estimate (Subsection 4.1) and the bagged 1-nearest neighbor estimate (Section 4.2). As we will see, both satisfy the deviation inequality (2).
@@ 4.2. Bagging and Nearest Neighbors
Here we focus on sampling with replacement and get the following bound, whose proof can be found in Section 6.3. The case of sampling without replacement leads to the same type of result.